OpenStudy (anonymous):

i/2-i how would i do this?

OpenStudy (anonymous):

it doesnt have it im looking for it in the book i have no idea

OpenStudy (anonymous):

i equals negative 1 the book

OpenStudy (anonymous):

sqyareoot of -1

OpenStudy (anonymous):

i got -0.5 i dnt thik thts correct

OpenStudy (anonymous):

okk il read it

OpenStudy (anonymous):

iv read them

OpenStudy (anonymous):

wasnt it the answer i gave u before ?

OpenStudy (anonymous):

\[ \frac{i}{2}-i = \frac{i}{2}-\frac{2i}{2} = -\frac{i}{2} \]

OpenStudy (anonymous):

-0.5 right ?

OpenStudy (anonymous):

The \(i\) should not disappear.

OpenStudy (anonymous):

0.5i then

OpenStudy (anonymous):

The negative should not disappear either.

OpenStudy (anonymous):

i mean -

OpenStudy (anonymous):

-0.5i

OpenStudy (anonymous):

lol so my answer is right right

OpenStudy (unklerhaukus):

\[\frac i{2-i}\\=\frac i{2-i}\times1\\=\frac i{2-i}\times\frac {2+i}{2+i}\\=\frac{2i+i^2}{4-i^2}\\=\frac{2i-1}{4+1}\\=\frac{2i-1}{5}\\=\tfrac25i-\tfrac15\]

OpenStudy (anonymous):

harddddd

OpenStudy (anonymous):

thank u guys :)

OpenStudy (anonymous):

@unklerhaukus the answer comes down to 0.2 i but i got 0.5i before